OpenStudy (anonymous):

A cooler contains 5 bottles of orange juice, 6 bottles of apple juice and 4 bottles of grape juice. Janice randomly selects a bottle from the cooler. Then David selects a bottle from the cooler. What is the probability that both Janice and David will select orange juice

OpenStudy (anonymous):

Joint probability is the product of the individual probabilities. Janice: p1 = 5/(5 + 6 + 4) and she uses up one OJ David: p2 = 4/(4 + 6 + 4) together p = p1 p2 = 20 / (15)(14)

OpenStudy (anonymous):

5+6+4=15. Orange juice is 5/15. Probability of Janice getting OJ is 5/15. 5-1=4 4+6+4=14 Remaining OJ=4 so the probability of David getting OJ is 4/14. You then add the two probabilities together and you get 13/21... I think.

OpenStudy (anonymous):

No. Do not add these. See my reply above. Multiply.